OpenStudy (ksaimouli):

plz help me (x+1)(x^2-3x+2)<0 slove the polynomial inequalities using sign chart plz help me

OpenStudy (ksaimouli):

help me bro

OpenStudy (anonymous):

(x+1) = 0 OR (x^2 -3x +2) =0 x = -1 OR X=1 OR x = 2 

OpenStudy (ksaimouli):

how did u get - for 1 and 2

OpenStudy (ksaimouli):

iam stuck their

OpenStudy (anonymous):

Just fill in a value in the function. between 1 and 2 choose 1.5 or smthng

OpenStudy (anonymous):

(x+1)(x^2-3x+2) x = 1.5 1.5 + 1 is positive 1.5^2 -4.5 +2 is negative, hence it's negative

OpenStudy (ksaimouli):

ok got u thx bro